38-Assume the price of good Y with its quantity measured on the measured on the horizontal axis is $5. If the consumers budget is $100, then the absolute value of the slope of the budget line is: vertical axis is $20 and the price of good X with its quantity a. 100. b. 20. 39-If the exchange rate between the yen and the dollar changes from 100 yen $1 to 110 yen $1, then: a. the dollar has depreciated in value. b. U.S.-made goods will become less expensive to Japanese citizens. c. the dollar has appreciated in value. d. Japanese-made goods will become more expensive to U.S. citizens. e. there will be an increase in the demand for dollars in the foreign exchange market. 40-Pauls Plumbing is a small business that employs 12 people. Which of the following is the best example of an implicit cost incurre by this firm? a. The tax payments on property owned by the firm. b. The wages paid to the 12 employees. c. The half of the payroll taxes on the wages of the 12 employees paid by the employers, but not the half paid by the employees. d. The accounting services provided free of charge to the firm by Pauls wife, who is an accountant. 41-Identify whether each of the following items creates a surplus item or a deficit item in the current account of the U.S balance of payments. a. A Central European company sells products to a U.S. hobby-store chain. b. Japanese residents pay a U.S. travel company to arrange hotel stays, ground transportation, and tours of various U.S cities, including New York, Chicago, and Orlando. c. A Mexican company pays a U.S. accounting firm to audit its income statements. d. U.S. churches and mosques send relief aid to Pakistan following a major earthquake in that nation. e. A U.S. microprocessor manufacturer purchases raw materials from a Canadian firm.
Answer #1

38. Option c. = Slope = 5/20 = 1/4

39. Option c. As US dollar can buy more Japanese goods

40. Option d. As is not paid out cost

41. a. As it falls under imports, it creates a deficit

b. As it falls under exports, it creates a surplus

c. As it falls under exports, it creates a surplus

d. It represents deficit

e. As it falls under imports, it creates a deficit